OpenStudy (anonymous):

solve the equation : 2^(4x-1)=18 ?

OpenStudy (anonymous):

ln(2^(4x-1) = ln(2*3^2) (4x-1) ln2= ln2 + 2ln3 4x-1=1 +2 (ln3/ln2) 4x=2+2 (ln3/ln2) x=1/2 +1/2 * (ln3/ln2)